OpenStudy (anonymous):

Is this the right answer. questions: use properties of logarithms to solve each equation 10^1-2x = 72 my answer 1.43

OpenStudy (anonymous):

log 10 - 2x log 10 = log 72 is that what you did?

OpenStudy (anonymous):

nope

OpenStudy (zehanz):

Property of logarithms:\[\log_{b}a=c <=>a=b^c \] This means that you can rewrite an equation like yours in an exponential form and vice versa:\[10^{1-2x}=72 <=>\log_{10}72=1-2x \]Now you can go on solving x...